OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? Which of the following sets is a Pythagorean triple? a. 5, 6,7 b. 6,7,8 c. 5,12, 13 d. 12, 13, 14

OpenStudy (anonymous):

*not a*

OpenStudy (anonymous):

C) 5,12,13

OpenStudy (paxpolaris):

keep trying, does \(a^2+b^2=c^2\) ... where \(c \) is the largest of the 3 numbers

OpenStudy (triciaal):

from practice we know its C to determine you would square the largest number if this is the same as the sum of the squares of the other 2 numbers then you have a pythagorean triple a^2 + b^2 = c^2 c is the largest number in the set

OpenStudy (triciaal):